HUBBLE SPACE TELESCOPE DAILY REPORT #1811 PERIOD COVERED: 0000Z (UTC) 02/06/97 - 0000Z (UTC) 02/07/97 Daily Status Report as of 038/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 6520 (The Origins of Boron) The High Resolution Spectrograph (Side One) was used to make = observations of BD-06D1250 (a star located in the Orion class OB1 = association) in order to study the origins of the element boron. This = star is the youngest of these stars and its shows higher oxygen = abundances than the other older stars. Approximately four-and-a-half = hours into this ten-and-a-half hour proposal, the GHRS safed while = operating with high voltage on Side 2. HSTAR 5889 was written to = describe this and the resulting status buffer messages. See Section 2.3 = for further details. 1.2 Completed Two Sets of WF/PC-2 6909 (WF/PC-2 Cycle 6 Calibration: = Earth Flats The WF/PC-2 was used to make observations of the bright Earth to = obtain streak flats which are used to construct high quality flat fields = for the WF/PC-2 filter set. The observations were executed as scheduled, = and no anomalies were noted. 1.3 Completed WF/PC-2 6906 (WF/PC-2 Cycle 6 Calibration: Internal = Monitor-Visual Flats) The WF/PC-2 was used to make internal calibration visual flat = images as part of the Cycle 6 internal monitoring of the instrument's = operations. This proposal defines the standard series of internal flat = fields, which will be executed twice per month. The observations were = completed as planned, and no anomalies were reported. 1.4 Completed FOS/BL 6584 (Interaction of Supernovae With = Curcumstellar Material) The Fine Object Spectrograph (Blue Side) was used to make = observations of Supernova SN1995N. The objective is to study the = interaction between the supernova and the remnant circumstellar material = from the progenitor star, and how that interaction can sustain supernovae = luminosity for long periods. Two hours before the end of this ten-and-a-h= alf hour proposal, the failed re-acquisition mentioned in 2.1 occurred. = One observation was taken with the take data flag down. Otherwise, the = remaining observations executed as scheduled, and no further anomalies = were noted. 1.5 Completed Two Sets of WF/PC-2 7203 (Cosmology with the Medium Deep = Survey - Short Orbit Cases) The WF/PC-2 was used to take short (about 20-minute) exposures, to = be taken in parallel with the prime science observations, using the I = filter. These exposures are part of the medium-deep survey, which = emphasizes exposures that will be useful in obtaining cosmological = parameters. The observations were completed as planned, and no problems = were reported. 1.6 Completed WF/PC-2 6802 (Cosmology with The Medium Deep Survey) The WF/PC-2 was used to make observations in parallel with the = prime science observations as part of the medium deep survey (MDS). = These parallel MDS exposures have provided unique high resolution images = which clearly supplement the primary HST surveys. During this proposal, = the failed re-acquisition mentioned in 2.1 occurred. One observation was = taken with the take data flag down. Otherwise, the remaining = observations executed as scheduled, and no further anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 11 Successful : 10 Per HSTAR 5890 and at 037/203801Z, a guide star re-acquisition = failed. 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Operations Notes: At 037/045901, GHRS safed due to limit violations on Z28AIN2, the = input current to Side 2. The PDU current showed a jump of 7 amps at the = same time. GHRS was in the middle of a series of observations, with the = HV on, but was not observing at the time. Numerous other GHRS voltage = monitors show step changes at the same time. No other unusual spacecraft = events were noted. At 037/2231Z, after a series of meetings, the LVPS = for Side 1 was turned on. The GHRS COSTAR arm is scheduled to be retracted on 040/0905Z in = preparation for the GHRS change-out for SM2. Other meetings today will = decide whether the COSTAR observations will occur as planned or whether a = workaround will be utilized. 3.0 SIGNIFICANT FORTHCOMING EVENTS: SM2 is scheduled to begin next Tuesday, with the launch of the = Shuttle at approximately 042/0856Z. /DMH